# Polynomial bounds for surfaces in cusped 3-manifolds

### v2

@@ 8.2. Orientable surfaces in thickened surfaces
When we restrict to orientable surfaces, more results are known. For example, the following theorem appears in a paper of Waldhausen from 1968 [39, Corollary 3.2].
@@
Let $\pi(L)$ be a weakly generalised alternating projection of a link $L$ onto a projection surface $\Pi=F\times\{0\}\subset Y=F\times[-1,1]$, such that $\pi(L)$ is cellular, and has $n$ crossings. Fix a connected, orientable topological surface $Z$ with Euler characteristic $\chi$. Then up to isotopy, the number of ways to properly embed $(Z,\partial Z)$ as an essential, meridianally incompressible surface in $(Y-N(L),\partial N(L))$ is at most:
@@
Note first that $Y$, $\Pi$, and $L$ satisfy Assumptions 2.1 and 2.9. Thus Theorem 7.1 and Theorem 7.4 will apply, once we determine a value for $X$. The chunk decomposition consists of two chunks of the form $F\times I$. Any embedded essential surface in $Y-N(L)$ with boundary on $N(L)$, when isotoped into normal form, will have boundary on only one side of the chunk, namely the boundary component of $F\times I$ that meets the diagram $\pi(L)$. Then Theorem 8.5 implies that such a surface is boundary parallel when orientable, hence uniquely determined by its boundary curves, with the surface lying on at most two sides of a component of the boundary curve. For these manifolds, we may set $X=2$. ∎
@@ 9. Dehn fillings
Let $\pi(L)$ be a weakly generalised alternating projection of a knot $L$ onto a projection surface $\Pi$ in a compact, irreducible, orientable 3-manifold $Y$, satisfying Assumptions 2.1 and 2.9. Let the number of crossings of $\pi(L)$ be $n$. Suppose that there is a universal bound $X$ on the number of isotopy classes of incompressible surfaces properly embedded in $\Sigma$ with fixed genus and a fixed finite set of boundary curves on $\partial N(\Pi)\cap\partial\Sigma$.
Finally, suppose $Y-L$ is hyperbolic. Take $(p,q)$ to be a nonmeridianal slope of length $\operatorname{len}(p,q)>2\pi$, and let $Q=4\pi(g-1)/(\operatorname{len}(p,q)-2\pi)$.
Then the number of isotopy classes of closed orientable genus $g$ surfaces in the manifold obtained by $(p,q)$ Dehn filling on $Y-L$ is at most
@@
Hass, Rubinstein, and Wang showed in [14, Theorem 4.1] that if $Y-L$ is hyperbolic, then for all but finitely many Dehn filling slopes, the essential surfaces of genus $g$ in the Dehn filling of $Y-L$ along that slope are exactly the essential surfaces in $Y-L$. Thus for these slopes, the upper bound comes from Theorem 7.4. In general, for all slopes, this is the first term in the sum.
@@
Let $S$ be an essential surface of genus $g$ embedded in $Y-N(L)$ with $b$ boundary components, each of slope $(p,q)$ on $\partial N(L)$. We may give $S$ a pleating (see, for example [35, Proposition 8.40]); $S$ then inherits a complete hyperbolic metric from the pleating in $Y-N(L)$. Note that the area of $S$ within the cusp of $Y-L$ is at least equal to $b$ times the length of the slope $(p,q)$ (e.g. [35, Lemma 8.44]). Thus $\operatorname{area}(S)\geq b\operatorname{len}(p,q)$.
On the other hand, by the Gauss–Bonnet formula we have $\operatorname{area}(S)=-2\pi\chi(S)=2\pi(2g+b-2)$. It follows that
@@
If we adjust our hypotheses on the link and the filling slope slightly, the fact that $Y-L$ is hyperbolic will be automatic, and similarly we obtain a bound on the slope length, giving the following.
@@
Let $\pi(L)$ be a weakly generalised alternating projection of a knot $L$ onto a projection surface $\Pi$, with $n$ crossings, in a 3-manifold $Y$, satisfying Assumptions 2.1 and 2.9 as usual. Suppose further that $Y-N(\Pi)$ contains no essential annuli with both boundary components on $\partial Y$. Suppose there is a universal bound $X$ on the number of isotopy classes of incompressible surfaces properly embedded in each component of $Y-N(\Pi)$ with fixed genus and fixed boundary. Let $\sigma=(p,q)$ be a slope on $\partial N(K)$, with $|q|>5.627(1-\chi(\Pi)/n)$.
Then the number of isotopy classes of closed orientable genus $g$ surfaces in the manifold obtained by $(p,q)$ Dehn filling of $Y-L$ is bounded above by a polynomial function of $n$, equal to the bound of (9.2).
@@
Under the given assumptions, $Y-L$ is hyperbolic by [19, Theorem 1.1]. It is shown in the proof of [19, Corollary 7.2] that the length of $\sigma$ is at least
@@
First suppose $F$ is both incompressible and boundary incompressible. If it is orientable, then it must be vertical or horizontal; see for example [20, Theorem VI.34]. In $T^{2}\times I$, a vertical surface is of the form $\{\gamma\}\times I$, and a horizontal surface is a torus parallel to the boundary. A non-orientable surface will be pseudo-horizontal or pseudo-vertical in general, as defined and proved by Frohman [10]. However, in $T^{2}\times I$, both of these definitions reduce to the above notion of horizontal and vertical, and we are done.
Now assume $F$ is boundary compressible.
If a component of $\partial F$ bounds a disk $D$ on $T^{2}\times\{0\}$ or $T^{2}\times\{1\}$, then because $F$ is incompressible, $\partial D$ pushed slightly into $F$ bounds a disk in $F$; it follows that $F$ is a boundary parallel disk, contradicting the hypothesis. So we may suppose that each component of $\partial F$ is essential.
@@
So suppose that the arc $\alpha$ has both endpoints on the same curve of $\partial F$. The arc $\beta$ cannot cut off a disk $E$ in $T^{2}\times\{0\}-\partial F$, else $E\cup D$ would be a compression disk for $F$. Thus $\beta$ is an essential arc in an annulus $T^{2}\times\{0\}-\partial F$. It follows that $\partial F$ has only one component on $T^{2}\times\{0\}$. Thus an arc $\gamma$ of $\partial F$ runs from one endpoint of $\partial\alpha$ to the other.
Because $\gamma$ is an essential curve on the torus, it runs from one side of $D$ to the opposite side, as in Figure 3, left. It cannot run to the same side of $D$, else $\gamma\cup\beta$ would bound a disk $E$ on $T^{2}\times\{0\}$, and $E\cup D$ would be a compressing disk for $F$.
@@
Now repeat the argument, applied to $F\cap T^{2}\times[\epsilon_{1},1]$. If there is another boundary compressing disk toward $T^{2}\times\{\epsilon_{1}\}$, then again there will be some $\epsilon_{2}>\epsilon_{1}$ so that $T^{2}\times[\epsilon_{1},\epsilon_{2}]\cap F$ is a Möbius band with a hole. Thus $F$ is obtained by stacking two of these, yielding a nonorientable surface with genus 22 22, and unique boundary components on $T^{2}\times\{0\}$ and $T^{2}\times\{\epsilon_{2}\}$, with the slope on $T^{2}\times\{\epsilon_{2}\}$ intersecting that on $T^{2}\times\{\epsilon_{1}\}$ exactly twice. This process must terminate with some $\epsilon_{n}$, because the Euler characteristic of $F$ is finite. This implies that $\epsilon_{n}=1$, and we have constructed a nonorientable surface with two boundary components on $T^{2}\times\{0\}$ and $T^{2}\times\{1\}$, as in case (3). This proves that any incompressible surface in $T^{2}\times I$ is as claimed.
@@
Finally, consider uniqueness of the nonorientable surfaces with fixed boundary components. At each step of the proof, we built a surface that is a Möbius band with a hole, with one boundary component on $T^{2}\times\{\epsilon_{i}\}$ and one on $T^{2}\times\{\epsilon_{i+1}\}$, and the boundary curves intersecting exactly twice. Adjust the framing on $T^{2}\times\{0\}$ so that the original slope $F\cap T^{2}\times\{0\}$ is $0/1$, and together with a choice of slope $1/0$ this forms a basis for the fundamental group of $T^{2}$. The construction above then starts with the slope $0/1$, and adds a band to the surface to obtain a new boundary slope of the form ±$\pm 2/q$, intersecting $0/1$ exactly twice. At each step, we replace a slope with even numerator by one meeting the first exactly twice, hence the result continues to have even numerator.
@@
Recall that slopes on the torus correspond to elements of ${\mathbb{Q}}\cup\{1/0\}$. These can be viewed as vertices of the Farey triangulation of ${\mathbb{H}}^{2}$: identifying ${\mathbb{H}}^{2}$ with the upper half plane, the vertices of the Farey triangulation are points of ${}\displaystyle{\mathbb{Q}}\cup\{\infty\}$ on the real line, and the edges run between reduced pairs $p/q$, $r/s$ if and only if $ps-qr=\pm 1$. Here, $ps-qr$ is the (signed) intersection number of the slopes.
@@
We consider slopes with even numerator and draw an edge between them if they have intersection number ±$\pm 2$. Equivalently, we may consider slopes with odd denominator and the edges between them with intersection number ±$\pm 1$. This gives a subset of the usual Farey triangulation. Observe that this subset, consisting only of vertices with odd denominators and the edges between them, forms a connected tree; see for example [4, Section 3.1]. Thus for any even slope $2p/q$, by following the unique path in this subset of the Farey tree from $0/1$ to $2p/q$, we may build a nonorientable surface in $T^{2}\times[0,1]$ with slopes $0/1$ on $T^{2}\times\{0\}$ and $2p/q$ on $T^{2}\times\{1\}$, where following an edge corresponds to adding an appropriate band to the surface.
Observe that traversing an edge, and then returning along it right away, constructs a surface with a compression disk. Thus the construction proceeds monotonically through the Farey tree, and the surface is unique up to isotopy, by uniqueness of the path through the Farey tree. Therefore the pair of slopes uniquely determines the nonorientable surface. ∎

### v3

@@ 8.2. Orientable surfaces in thickened surfaces
When we restrict to orientable surfaces, more results are known. For example, the following theorem appears in a paper of Waldhausen from 1968 [54, Corollary 3.2].
@@
Theorem 8.5 allows us to apply a bound to cellular weakly generalized alternating links with projection surface $\Pi=F\times\{0\}$ in a thickened surface $Y=F\times[-1,1]$. These are virtual alternating links. Virtual links were introduced by Kauffman [35], and subsequently shown to be equivalent to diagrams of links in thickened surfaces up to moves called stabilisation and destabilisation. Kuperberg proved there exists a unique minimal genus surface $F$ for which the link embeds in $F\times[-1,1]$ with diagram on $F\times\{0\}$ admitting no destabilisations [37]. Such a diagram must be cellular. If we require the diagram to be weakly prime and alternating, it is exactly a weakly generalized alternating link, and satisfies all assumptions required for our theorems.
@@
Let $\pi(L)$ be a virtual link $L$ with a weakly prime alternating projection onto a surface $\Pi=F\times\{0\}\subset Y=F\times[-1,1]$ with $n$ crossings. Fix a connected, orientable topological surface $Z$ with Euler characteristic $\chi$. Then up to isotopy, the number of ways to properly embed $(Z,\partial Z)$ as an essential, meridianally incompressible surface in $(Y-N(L),\partial N(L))$ is at most:
@@
If in addition $Y-L$ admits a hyperbolic structure, then the above bounds quasifuchsian surfaces.
@@
Note first that $L$ is a weakly generalized alternating link, with infinite representativity because $\Pi$ has no compression discs in $Y$. Note also that $Y$, $\Pi$, and $L$ satisfy Assumptions 2.1 and 2.9. Thus Theorem 7.1 and Theorem 7.5 will apply, once we determine a value for $X$. The chunk decomposition consists of two chunks of the form $F\times I$. Any embedded essential surface in $Y-N(L)$ with boundary on $N(L)$, when isotoped into normal form, will have boundary on only one side of the chunk, namely the boundary component of $F\times I$ that meets the diagram $\pi(L)$. Then Theorem 8.5 implies that such a surface is boundary parallel when orientable, hence uniquely determined by its boundary curves, with the surface lying on at most two sides of a component of the boundary curve. For these manifolds, we may set $X=2$. ∎
@@ 9. Dehn fillings
Let $\pi(L)$ be a weakly generalized alternating projection of a knot $L$ onto a projection surface $\Pi$ in a compact, irreducible, orientable 3-manifold $Y$, satisfying Assumptions 2.1 and 2.9. Let the number of crossings of $\pi(L)$ be $n$. Suppose that there is a universal bound $X$ on the number of isotopy classes of incompressible surfaces properly embedded in $\Sigma$ with fixed genus and a fixed finite set of boundary curves on $\partial N(\Pi)\cap\partial\Sigma$. Finally, suppose $Y-L$ is hyperbolic. Take $(p,q)$ to be a nonmeridianal slope of length $\operatorname{len}(p,q)>2\pi$, and let $Q=4\pi(g-1)/(\operatorname{len}(p,q)-2\pi)$. Then the number of isotopy classes of closed orientable genus $g$ surfaces in the manifold obtained by $(p,q)$ Dehn filling on $Y-L$ is at most
@@
Hass, Rubinstein, and Wang showed in [23, Theorem 4.1] that if $Y-L$ is hyperbolic, then for all but finitely many Dehn filling slopes, the essential surfaces of genus $g$ in the Dehn filling of $Y-L$ along that slope are exactly the essential surfaces in $Y-L$. Thus for these slopes, the upper bound comes from Theorem 7.5. In general, for all slopes, this is the first term in the sum.
@@
Let $S$ be an essential surface of genus $g$ embedded in $Y-N(L)$ with $b$ boundary components, each of slope $(p,q)$ on $\partial N(L)$. We may give $S$ a pleating (see, for example [48, Proposition 8.40]); $S$ then inherits a complete hyperbolic metric from the pleating in $Y-N(L)$. Note that the area of $S$ within the cusp of $Y-L$ is at least equal to $b$ times the length of the slope $(p,q)$ (e.g. [48, Lemma 8.44]). Thus $\operatorname{area}(S)\geq b\operatorname{len}(p,q)$.
By Gauss–Bonnet, we also have $\operatorname{area}(S)=-2\pi\chi(S)=2\pi(2g+b-2)$. It follows that
@@
If we adjust our hypotheses on the link and the filling slope slightly, $Y-L$ will automatically be hyperbolic with a bound on the slope length.
@@
Let $\pi(L)$ be a weakly generalized alternating projection of a knot $L$ onto a projection surface $\Pi$, with $n$ crossings, in a 3-manifold $Y$, satisfying Assumptions 2.1 and 2.9. Suppose that $Y-N(\Pi)$ contains no essential annuli with both boundary components on $\partial Y$. Suppose there is a universal bound $X$ on the number of isotopy classes of incompressible surfaces properly embedded in each component of $Y-N(\Pi)$ with fixed genus and fixed boundary. Let $\sigma=(p,q)$ be a slope on $\partial N(K)$, with $|q|>5.627(1-\chi(\Pi)/n)$. Then the number of isotopy classes of closed orientable genus $g$ surfaces in the manifold obtained by $(p,q)$ Dehn filling of $Y-L$ is bounded above by a polynomial function of $n$, equal to the bound of (9.2).
@@
Under the given assumptions, $Y-L$ is hyperbolic by [27, Theorem 1.1]. It is shown in the proof of [27, Corollary 7.2] that the length of $\sigma$ is at least
@@
First suppose $F$ is both incompressible and boundary incompressible. If it is orientable, then it must be vertical or horizontal; see for example [29, Theorem VI.34]. In $T^{2}\times I$, a vertical surface is of the form $\{\gamma\}\times I$, and a horizontal surface is a torus parallel to the boundary. A non-orientable surface will be pseudo-horizontal or pseudo-vertical in general, as defined and proved by Frohman [19]. However, in $T^{2}\times I$, both of these definitions reduce to the above notion of horizontal and vertical.
Now assume $F$ is boundary compressible. If a component of $\partial F$ bounds a disk $D$ on $T^{2}\times\{0\}$ or $T^{2}\times\{1\}$, then because $F$ is incompressible, $\partial D$ pushed slightly into $F$ bounds a disk in $F$; it follows that $F$ is a boundary parallel disk, contradicting the hypothesis. So we may suppose that each component of $\partial F$ is essential.
@@
So suppose that the arc $\alpha$ has both endpoints on the same curve of $\partial F$. The arc $\beta$ cannot cut off a disk $E$ in $T^{2}\times\{0\}-\partial F$, else $E\cup D$ would be a compression disk for $F$. Thus $\beta$ is an essential arc in an annulus $T^{2}\times\{0\}-\partial F$. It follows that $\partial F$ has only one component on $T^{2}\times\{0\}$. Thus an arc $\gamma$ of $\partial F$ runs from one endpoint of $\partial\alpha$ to the other. Because $\gamma$ is an essential curve on the torus, it runs from one side of $D$ to the opposite side, as in Figure 3, left. It cannot run to the same side of $D$, else $\gamma\cup\beta$ would bound a disk $E$ on $T^{2}\times\{0\}$, and $E\cup D$ would be a compressing disk for $F$.
$\beta$
$D$
$T^{2}\times\{0\}$
$T^{2}\times\{\epsilon_{1}\}$
$N(\alpha)$
$N(\partial F)$
@@
Repeat the argument, applied to $F\cap T^{2}\times[\epsilon_{1},1]$. If there is another boundary compressing disk toward $T^{2}\times\{\epsilon_{1}\}$, then there will be some $\epsilon_{2}>\epsilon_{1}$ so that $T^{2}\times[\epsilon_{1},\epsilon_{2}]\cap F$ is a Möbius band with a hole. Thus $F$ is obtained by stacking two of these, yielding a nonorientable surface with genus $2$, and unique boundary components on $T^{2}\times\{0\}$ and $T^{2}\times\{\epsilon_{2}\}$, with the slope on $T^{2}\times\{\epsilon_{2}\}$ intersecting that on $T^{2}\times\{\epsilon_{1}\}$ exactly twice. This process must terminate with some $\epsilon_{n}$, because the Euler characteristic of $F$ is finite. This implies that $\epsilon_{n}=1$, and we have constructed a nonorientable surface with two boundary components on $T^{2}\times\{0\}$ and $T^{2}\times\{1\}$, as in case (3).
@@
Finally, consider uniqueness of the nonorientable surfaces with fixed boundary components. At each step of the proof, we built a surface that is a Möbius band with a hole, with one boundary component on $T^{2}\times\{\epsilon_{i}\}$ and one on $T^{2}\times\{\epsilon_{i+1}\}$, and the boundary curves intersecting exactly twice. Adjust the framing on $T^{2}\times\{0\}$ so that the original slope $F\cap T^{2}\times\{0\}$ is $0/1$, and together with a choice of slope $1/0$ this forms a basis for the fundamental group of $T^{2}$. The construction above then starts with the slope $0/1$, and adds a band to the surface to obtain a new boundary slope of the form $\pm 2/q$, intersecting $0/1$ exactly twice. At each step, we replace a slope with even numerator by one meeting the first exactly twice, hence the result continues to have even numerator.
@@
Recall that slopes on the torus correspond to elements of ${\mathbb{Q}}\cup\{1/0\}$. These can be viewed as vertices of the Farey triangulation of ${\mathbb{H}}^{2}$: identifying ${\mathbb{H}}^{2}$ with the upper half plane, the vertices of the Farey triangulation are points of ${\mathbb{Q}}\cup\{\infty\}$ on the real line, and the edges run between reduced pairs $p/q$, $r/s$ if and only if $ps-qr=\pm 1$. Here, $ps-qr$ is the (signed) intersection number of the slopes.
@@
We consider slopes with even numerator and draw an edge between them if they have intersection number $\pm 2$. Equivalently, we may consider slopes with odd denominator and the edges between them with intersection number $\pm 1$. This gives a subset of the usual Farey triangulation. Observe that this subset, consisting only of vertices with odd denominators and the edges between them, forms a connected tree; see for example [4, Section 3.1]. Thus for any even slope $2p/q$, by following the unique path in this subset of the Farey tree from $0/1$ to $2p/q$, we may build a nonorientable surface in $T^{2}\times[0,1]$ with slopes $0/1$ on $T^{2}\times\{0\}$ and $2p/q$ on $T^{2}\times\{1\}$, where following an edge corresponds to adding an appropriate band to the surface. Observe that traversing an edge, and then returning along it right away, constructs a surface with a compression disk. Thus the construction proceeds monotonically through the Farey tree, and the surface is unique up to isotopy, by uniqueness of the path through the Farey tree. Therefore the pair of slopes uniquely determines the nonorientable surface. ∎